# When Bias Helps Learning: Bridging Initial Prejudice and Trainability

### v3

@@ B.3 Vanilla Vision Transformer
We employ a vanilla Vision Transformer (VIT) [dosovitskiy2020image] for classification trained on binarized CIFAR10 (Fig. 11) and CIFAR10 (Fig. 12). We eliminate all batch- and layer-norm layers, since they are not included in our theory and could potentially destroy the phase distinctions. We preserve only linear and convolutional layers, which we initialize with the prescription of this work. In particular, this assumption is justified since the MF theory of convolutional layer is identical to that of MLP [pmlr-v80-xiao18a]. We initialize the attention matrices as the linear layer, with the gain depending on the activation function placed before the activation. This way, we show that gradients manifest the same transition behaviour observed in vanilla models, where an exact theory is available (see for instance Fig. 24).
@@ B.4 Large Vision Transformer
Finally, we validate our assumption by fine-tuning a large model, without any architectural modification. We download a pre-train large Vision Transformer (vit_large_patch16_384) from the library Timm [rw2019timm]. This model was pre-trained on ImagenNet-21k, and consists of about $300$ million parameters. We fine tune this model on CIFAR100. We multiply the weights of all linear and convolutional layers by a factor $\sigma^{2}_{w}$, with the goal of triggering the order/chaos phase transition studied in this paper. This way, the pre-trained model behaves like an untrained model when evaluated on a statistically different dataset. Indeed, we show that gradients behave like at the phase boundary between order and chaos (Fig. 25). We report the training dynamics in Fig. 13, where we observe that the original unscaled state (at EOC - red line) manifests the fastest learning dynamics. In particular, it begins with moderate prejudice and it is the only state significantly improving the classification accuracy of the most unfavoured class. Instead, the state rescaled with $\sigma^{2}_{w}=1.5$, ideally in the chaotic phase, begins with a lower amount of prejudice, but it cannot recover. Finally, the state rescaled by $\sigma^{2}_{w}=0.5$ shows strong initialization prejudice, typical of the ordered phase, but smaller gradients (Fig. 25), which hinder training.
@@ Appendix C Mean field results
In this section, we report some results of previous MF theory, starting from wide MLPs [poole2016exponential, schoenholz2016deep, hayou2019impact].
@@
Let us denote with $q$ the limiting variance in its domain of convergence [hayou2019impact]. We can show [schoenholz2016deep] that:
@@
In previous MF works, it was not clear the role played by $\chi_{1}$, because it is usually derived by assuming the variance to converge faster than the correlation coefficient [poole2016exponential, schoenholz2016deep]. Here, we prove that when the variance is not assumed to be constant, the result slightly changes, suggesting that for some activation functions, the asymptotic correlation coefficient is not able to discriminate between phases. First, we prove the following Lemma.
@@
Let us compute $\frac{\partial c_{ab}^{(l+1)}}{\partial c_{ab}^{(l)}}$. For generic activation functions, $q_{aa}^{(l)}$ may diverge with depth and thus it cannot be safely kept constant as done in the MF literature for bounded activation functions [schoenholz2016deep]. Moreover, due to our main result (reported as Thm. E.1), for large $N$, $q_{aa}^{(l)}=q_{bb}^{(l)}$, $\forall a,b\in\mathcal{D},\forall l$; thus, we can write $q_{aa}^{(l)}\equiv q^{(l)}$, $\forall a\in\mathcal{D}$. From Eq. (8) we can directly calculate:
@@
Next, we proceed using the following key identity known as Stein’s lemma [stein1956inadmissibility]:
@@ Appendix D IGB extension to explicit initialization biases
Ref. [pmlr-v235-francazi24a] derived the pre-activation distributions of a MLP processing a random dataset in case of zero explicit initialization biases (Thm D.2, Appendix). Here, we generalize it to accomplish non-zero initialization biases.
@@
The result in Thm D.2 extends directly from prior work that considered the same setting but with zero bias terms. In particular, Ref. [pmlr-v235-francazi24a] proved that under the assumption of i.i.d. Gaussian data, fixed weights and large layer, the pre-activations
@@
For a general MLP (Eq. (LABEL:MLP_recursion)), the signal variance and the centers variance satisfy the following recursive equations: $\sigma^{2}_{y^{(l+1)}}$ $=\sigma^{2}_{w}\textrm{Var}_{\mathcal{D}}\left({\mathrm{\phi}\left({y^{(l)}}\right)}\right)\;,$ (36) $\sigma^{2}_{\mu^{(l+1)}}$ $=\sigma^{2}_{w}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\mathrm{\phi}\left({y^{(l)}}\right)}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu+\sigma^{2}_{b}\;,$ (37) with initial values $\sigma^{2}_{y^{(0)}}=1$ $\sigma^{2}_{\mu^{(0)}}=0$. Moreover, $\textrm{Var}_{\mathcal{D}}\left({\mathrm{\phi}\left({y^{(l)}}\right)}\right)$ is self-averaging with respect to $\mathcal{W}$, that is $\textrm{Var}_{\mathcal{D}}\left({\mathrm{\phi}\left({y^{(l)}}\right)}\right)=\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\mathcal{D}}\left({\mathrm{\phi}\left({y^{(l)}}\right)}\right)\mkern-1.5mu}\mkern 1.5mu$ .
@@
Ref. [pmlr-v235-francazi24a] proved that in the large width limit $\textrm{Var}_{\mathcal{D}}\left({y_{i}^{(l)}}\right)$ is self-averaging (as distribution of the weights) and does not depend on $i$, i.e.
@@
Now let us consider the calculation for $\sigma^{2}_{\mu^{(l)}}$. From Eq. (LABEL:MLP_recursion) we easily see that $\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{y_{i}^{(l+1)}}\right\rangle\mkern-1.5mu}\mkern 1.5mu=0$ and
@@ Appendix E The equivalence between MF and IGB in the large-width limit
We plot the absolute percentage error for the ReLU of $\gamma^{(l)}$, $q_{aa}^{(l)}$, and $q_{ab}^{(l)}$as the network sizes increases in Figs. 14, 15, and 16, respectively. Moreover, we report the absolute percentage error of $q_{aa}^{(l)}$, and $q_{ab}^{(l)}$ for Tanh activation (Figs. 17, Fig. 18, respectively).
@@
Thm. E.1 establishes the equivalence between MF and IGB in the infinte-width limit. In MF, optimal training conditions are to be found at the edge of chaos (EOC). In the IGB framework, the EOC corresponds to the deep prejudice state, since the asymptotic value of the correlation coefficient is one. Moreover, this prejudiced state is transient since at the EOC, gradients are stable and so the network can absorb the bias rapidly [hayou2019impact]. ∎
@@ Appendix F MF/IGB extension to multi-node activation functions: the effect of pooling layers
Within the MF literature, common architectural components such as batch normalization [yang2019meanfieldtheorybatch] and dropout [schoenholz2016deep] have been extensively studied. In this work, however, we investigate the impact of multi-node pooling layers on the phase diagrams — a topic that has received comparatively little attention.
To this end, we leverage the IGB framework, which naturally extends to the general case of multi-node activation functions. We start by deriving recursive equations for a generic activation function $\phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$, involving $n$ nodes. Once these general recursive relations are established, the theoretical results from [hayou2019impact] remain applicable. In particular, we can directly employ Algorithm 1 from [hayou2019impact] to compute the EOC.
@@
For a generic function multi variable function $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$, [^1] In this proof, we omit the $l$-dependency for better readability. we have
@@ F.1 MaxPool
Let $f:\mathbb{R}^{2}\to\mathbb{R}^{2}$ a 2-node activation function that can be written as the composition of $\mathrm{MaxPool}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ with a single-node activation function $\phi:\mathbb{R}\to\mathbb{R}$. Then $\phi$ satisfies the following conditions 1. All conditions of Proposition 2 of the main paper of [hayou2019impact] 2. $\mathrm{\phi}\left({x}\right)$ is either odd or even 3. $\mathrm{\phi^{\prime}}\left({x}\right)$ is either odd or even Then $f=\mathrm{MaxPool}\circ\phi$ exhibit the same EOC of $\mathrm{\phi}\left({x}\right)$.
@@
From condition 1, we can use algorithm Algorithm 1 of [hayou2019impact] to compute the EOC. Then $f(\mathbf{x})$ exhibits the same EOC of $\mathrm{\phi}\left({x}\right)$ if and only if $V[f^{2}](x)=V[\phi^{2}](x)$ and $V[f^{\prime 2}](x)=V[\phi^{\prime 2}](x)$, $\forall x\in\mathbb{R}$. The $V$ operator is defined in Def. F.4. It is immediate to verify that condition 1 implies $V[f^{2}](x)=V[\phi^{2}](x)$ and condition 2 implies $V[f^{\prime 2}](x)=V[\phi^{\prime 2}](x)$. ∎
@@ F.2 Average Pooling
For Tanh, we can directly use Algorithm 1 of ref [hayou2019impact]. For ReLU, we have the following Lemma.
@@ Appendix G Gradients at initialization
Let be $\mathcal{L}$ the generic loss we want to optimize. The gradient compute for a data sample $a$ obeys the following equations:
@@
By assuming the forward weights to be independent from the backward ones, [hayou2019impact] proved that (see their Supplementary Materials)
@@ G.2 Residual MLP
We report the initialization gradients for a Residual MLP [NIPS2017_81c650ca] for both ReLU and Tanh activations on binarized CIFAR10 (BCIFAR - Fig. 23) For ReLU (Fig. 23 - top) we observe perfectly constant gradients, while for Tanh ( Fig. 23 - bottom) we observe a small exponential decay, which however does not hinder training and differentiate across phases (Figs. 9 and 10).
@@ G.3 Vanilla Vision Transformer
We compute the initialization gradients of a vanilla Vision Transformer [dosovitskiy2020image] with ReLU activation function on binarized CIFAR10 (BCIFAR, Fig. 24 - top) and CIFAR10 (Fig. 24 - bottom). By removing all batch- and layer-norm layers, we trigger the phase transition observed for instance in the MLP.
@@ G.4 Large Vision Transformer
Finally, we compute the gradients on CIFAR100 of a large Vision Tranformer pre-trained on Imagenet. The weights are rescaled as explained in Sec. B.4, while biases are not modified. We observe a similar transition behaviour as in vanilla models (Fig. 24. Due to computational limits, we are able to compute the full gradients only on a batch of size $100$ for this model.
@@
Appendix H Example of an explicit IGB calculations for the ReLU

### v4

@@ B.3 Vanilla Vision Transformer
We employ a vanilla Vision Transformer (VIT) (dosovitskiy2020image) for classification trained on binarized CIFAR10 (Fig. 12) and CIFAR10 (Fig. 13). We eliminate all batch- and layer-norm layers, since they are not included in our theory and could potentially destroy the phase distinctions. We preserve only linear and convolutional layers, which we initialize with the prescription of this work. In particular, this assumption is justified since the MF theory of convolutional layer is identical to that of MLP (pmlr-v80-xiao18a). We initialize the attention matrices as the linear layer, with the gain depending on the activation function placed before the activation. This way, we show that gradients manifest the same transition behaviour observed in vanilla models, where an exact theory is available (see for instance Fig. 27).
@@ B.4 Large Vision Transformer
Finally, we validate our assumption by fine-tuning a large model, without any architectural modification. We download a pre-train large Vision Transformer (vit_large_patch16_384) from the library Timm (rw2019timm). This model was pre-trained on ImagenNet-21k, and consists of about $300$ million parameters. We fine tune this model on CIFAR100. We multiply the weights of all linear and convolutional layers by a factor $\sigma^{2}_{w}$, with the goal of triggering the order/chaos phase transition studied in this paper. This way, the pre-trained model behaves like an untrained model when evaluated on a statistically different dataset. Indeed, we show that gradients behave like at the phase boundary between order and chaos (Fig. 28). We report the training dynamics in Fig. 14, where we observe that the original unscaled state (at EOC - red line) manifests the fastest learning dynamics. In particular, it begins with moderate prejudice and it is the only state significantly improving the classification accuracy of the most unfavoured class. Instead, the state rescaled with $\sigma^{2}_{w}=1.5$, ideally in the chaotic phase, begins with a lower amount of prejudice, but it cannot recover. Finally, the state rescaled by $\sigma^{2}_{w}=0.5$ shows strong initialization prejudice, typical of the ordered phase, but smaller gradients (Fig. 28), which hinder training.
@@ Appendix C Mean field results
In this section, we report some results of previous MF theory, starting from wide MLPs (poole2016exponential; schoenholz2016deep; hayou2019impact).
@@
Let us denote with $q$ the limiting variance in its domain of convergence (hayou2019impact). We can show (schoenholz2016deep) that:
@@
In previous MF works, it was not clear the role played by $\chi_{1}$, because it is usually derived by assuming the variance to converge faster than the correlation coefficient (poole2016exponential; schoenholz2016deep). Here, we prove that when the variance is not assumed to be constant, the result slightly changes, suggesting that for some activation functions, the asymptotic correlation coefficient is not able to discriminate between phases. First, we prove the following Lemma.
@@
Let us compute $\frac{\partial c_{ab}^{(l+1)}}{\partial c_{ab}^{(l)}}$. For generic activation functions, $q_{aa}^{(l)}$ may diverge with depth and thus it cannot be safely kept constant as done in the MF literature for bounded activation functions (schoenholz2016deep). Moreover, due to our main result (reported as Thm. E.1), for large $N$, $q_{aa}^{(l)}=q_{bb}^{(l)}$, $\forall a,b\in\mathcal{D},\forall l$; thus, we can write $q_{aa}^{(l)}\equiv q^{(l)}$, $\forall a\in\mathcal{D}$. From Eq. (8) we can directly calculate:
@@
Next, we proceed using the following key identity known as Stein’s lemma (stein1956inadmissibility):
@@ Appendix D IGB extension to explicit initialization biases
Ref. pmlr-v235-francazi24a derived the pre-activation distributions of a MLP processing a random dataset in case of zero explicit initialization biases (Thm D.2, Appendix). Here, we generalize it to accomplish non-zero initialization biases.
@@
The result in Thm D.2 extends directly from prior work that considered the same setting but with zero bias terms. In particular, Ref. pmlr-v235-francazi24a proved that under the assumption of i.i.d. Gaussian data, fixed weights and large layer, the pre-activations
@@
For a general MLP (Eq. (11)), the signal variance and the centers variance satisfy the following recursive equations: $\sigma^{2}_{y^{(l+1)}}$ $=\sigma^{2}_{w}\textrm{Var}_{\mathcal{D}}\left({\mathrm{\phi}\left({y^{(l)}}\right)}\right)\;,$ (36) $\sigma^{2}_{\mu^{(l+1)}}$ $=\sigma^{2}_{w}\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\mathrm{\phi}\left({y^{(l)}}\right)}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu+\sigma^{2}_{b}\;,$ (37) with initial values $\sigma^{2}_{y^{(0)}}=1$ $\sigma^{2}_{\mu^{(0)}}=0$. Moreover, $\textrm{Var}_{\mathcal{D}}\left({\mathrm{\phi}\left({y^{(l)}}\right)}\right)$ is self-averaging with respect to $\mathcal{W}$, that is $\textrm{Var}_{\mathcal{D}}\left({\mathrm{\phi}\left({y^{(l)}}\right)}\right)=\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\mathcal{D}}\left({\mathrm{\phi}\left({y^{(l)}}\right)}\right)\mkern-1.5mu}\mkern 1.5mu$ .
@@
Ref. pmlr-v235-francazi24a proved that in the large width limit $\textrm{Var}_{\mathcal{D}}\left({y_{i}^{(l)}}\right)$ is self-averaging (as distribution of the weights) and does not depend on $i$, i.e.
@@
Now let us consider the calculation for $\sigma^{2}_{\mu^{(l)}}$. From Eq. (11) we easily see that $\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{y_{i}^{(l+1)}}\right\rangle\mkern-1.5mu}\mkern 1.5mu=0$ and
@@ Appendix E The equivalence between MF and IGB in the large-width limit
We plot the absolute percentage error for the ReLU of $\gamma^{(l)}$, $q_{aa}^{(l)}$, and $q_{ab}^{(l)}$as the network sizes increases in Figs. 15, 16, and 17, respectively. Moreover, we report the absolute percentage error of $q_{aa}^{(l)}$, and $q_{ab}^{(l)}$ for Tanh activation (Figs. 18, Fig. 19, respectively).
@@
Thm. E.1 establishes the equivalence between MF and IGB in the infinte-width limit. In MF, optimal training conditions are to be found at the edge of chaos (EOC). In the IGB framework, the EOC corresponds to the deep prejudice state, since the asymptotic value of the correlation coefficient is one. Moreover, this prejudiced state is transient since at the EOC, gradients are stable and so the network can absorb the bias rapidly hayou2019impact. ∎
@@
E.1 MF/IGB equivalence for production architectures
To validate our theory in further realistic settings, we compute the correlation coefficient with both the IGB and MF theories for two production architectures downloaded from the Timm Python library: a RESNET with 18 layers (https://huggingface.co/docs/timm/en/models/resnet - Fig. 20) and a MLP mixer with a 100 layers (https://huggingface.co/timm/mixer_b16_224.miil_in21k_ft_in1k - Fig. 21). These figures should be interpreted in the same way as Fig. 2: the equivalence between the two frameworks is demonstrated by the fact that $c^{l}_{ab}$ has the same value independently of whether it is calculated through using the MF [Eq. (2)] or the IGB approach [Eq. (6)]. We employ the CIFAR10 dataset for these experiments.
Figure 20: Behaviour the correlation coefficient of a production RESNET. Solid lines are computed using the IGB approach, while shaded areas represent the 90 $\%$ central confidence interval computed using the mean-field approach. The IGB solid lines are all within the corresponding MF interval, demonstrating the agreement between the two theories. Moreover, coherently with residual MLPs (see the Appendix B.2), we do not observe distinction in the predicted correlation coefficient at varying $\sigma^{2}_{w}$.
Figure 21: Same as Fig. 20 for a production MLP-MIXER with 100 layers at initialization. The agreement between IGB and MF is good across the entire phase diagram, being stronger as we increase the scaling factor of the weights $\sigma^{2}_{w}$.
@@ Appendix F MF/IGB extension to multi-node activation functions: the effect of pooling layers
Within the MF literature, common architectural components such as batch normalization yang2019meanfieldtheorybatch and dropout schoenholz2016deep have been extensively studied. In this work, however, we investigate the impact of multi-node pooling layers on the phase diagrams — a topic that has received comparatively little attention.
To this end, we leverage the IGB framework, which naturally extends to the general case of multi-node activation functions. We start by deriving recursive equations for a generic activation function $\phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$, involving $n$ nodes. Once these general recursive relations are established, the theoretical results from (hayou2019impact) remain applicable. In particular, we can directly employ Algorithm 1 from (hayou2019impact) to compute the EOC.
@@
For a generic function multi variable function $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$,In this proof, we omit the $l$-dependency for better readability. we have
@@ F.1 MaxPool
Let $f:\mathbb{R}^{2}\to\mathbb{R}^{2}$ a 2-node activation function that can be written as the composition of $\mathrm{MaxPool}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ with a single-node activation function $\phi:\mathbb{R}\to\mathbb{R}$. Then $\phi$ satisfies the following conditions 1. All conditions of Proposition 2 of the main paper of hayou2019impact 2. $\mathrm{\phi}\left({x}\right)$ is either odd or even 3. $\mathrm{\phi^{\prime}}\left({x}\right)$ is either odd or even Then $f=\mathrm{MaxPool}\circ\phi$ exhibit the same EOC of $\mathrm{\phi}\left({x}\right)$.
@@
From condition 1, we can use algorithm Algorithm 1 of hayou2019impact to compute the EOC. Then $f(\mathbf{x})$ exhibits the same EOC of $\mathrm{\phi}\left({x}\right)$ if and only if $V[f^{2}](x)=V[\phi^{2}](x)$ and $V[f^{\prime 2}](x)=V[\phi^{\prime 2}](x)$, $\forall x\in\mathbb{R}$. The $V$ operator is defined in Def. F.4. It is immediate to verify that condition 1 implies $V[f^{2}](x)=V[\phi^{2}](x)$ and condition 2 implies $V[f^{\prime 2}](x)=V[\phi^{\prime 2}](x)$. ∎
@@ F.2 Average Pooling
For Tanh, we can directly use Algorithm 1 of ref hayou2019impact. For ReLU, we have the following Lemma.
@@ Appendix G Gradients at initialization
Let $\mathcal{L}$ be the generic loss we want to optimize. The gradient compute for a data sample $a$ obeys the following equations:
@@
By assuming the forward weights to be independent from the backward ones, hayou2019impact proved that (see their Supplementary Materials)
@@ G.2 Residual MLP
We report the initialization gradients for a Residual MLP (NIPS2017_81c650ca) for both ReLU and Tanh activations on binarized CIFAR10 (BCIFAR - Fig. 26) For ReLU (Fig. 26, top), the gradients remain perfectly constant across layers. For Tanh (Fig. 26, bottom), we observe a mild exponential decay—noticeably smaller than in the non-residual MLP (Fig. 25). However, this decay does not impede training nor the model’s ability to separate phases (Figs. 10 and 11).
@@ G.3 Vanilla Vision Transformer
We compute the initialization gradients of a vanilla Vision Transformer (dosovitskiy2020image) with ReLU activation function on binarized CIFAR10 (BCIFAR, Fig. 27 - top) and CIFAR10 (Fig. 27 - bottom). By removing all batch- and layer-norm layers, we trigger the phase transition observed for instance in the MLP.
@@ G.4 Large Vision Transformer
Finally, we compute the gradients on CIFAR100 of a large Vision Tranformer pre-trained on Imagenet. The weights are rescaled as explained in Sec. B.4, while biases are not modified. We observe a similar transition behaviour as in vanilla models (Fig. 27). Due to computational limits, we are able to compute the full gradients only on a batch of size $100$ for this model.
@@
G.4.1 The effect of the pre-training prior on initialization gradients
One question that might arise naturally is whether the phenomenology analyzed in this paper appears in the same way in pretrained and untrained models. We address this question by comparing the propagation of gradients in a pretrained vs an untrained ViT.
We compare the rescaled gradients of the pretrained large ViT with those of same model –this time untrained–, both computed on a batch of CIFAR100 (Fig. 29). Each hidden layer is organized as a block composed of multiple units (or elements). The regular oscillations observed within each layer reveal both the blockwise modular structure and the contribution of individual elements within each block (Fig. 29–b). To more clearly illustrate propagation effects across layers, we additionally report the gradient norms averaged over each layer block (Fig. 29–a). This aggregation suppresses within-layer oscillations and makes the inter-layer propagation behavior more apparent.
As shown in Sec. G.2, we expect that the presence of residual connections suppresses the gradient instability, putting the models closer to an EOC situation. This is indeed what we observe in Fig. 29–b. Surprisingly, however, we do observe gradient exploding in the pretrained netowrk (Fig. 29–a). This suggests that the prior induced by pretrained weights makes gradient stability substantially more fragile than in the untrained case, leading to noticeably stronger gradient explosion under the same range of $\sigma^{2}_{w}$.
Figure 29: Initialization gradients of the large VIT on CIFAR100, pre-trained and untrained models. In (a) we average the gradients over blocks of four layers to improve clarity, while the original gradients are reported in (b). For easier comparison, curves related to a different value of $\sigma^{2}_{w}$ are shifted in such a way to be equal to 1 at the last layer.
Appendix H Example of an explicit IGB calculation for the ReLU